OpenStudy (anonymous):

What is the equation of a line with a slope of -3/2 and x-intercept of (4,0)

OpenStudy (anonymous):

y - y1 = m (x-x1) (x1,y1) the point you are given m = slope plug the values in and re write into y = mx +b